OpenStudy (anonymous):

What is the solution for this system of linear equations: Tell me what I did wrong? :/ y=4x+13 y=-3x+6 SO, -3x+6=4x+13 subtracting 3 from both sides 6=1x+13 subtract 13 from both sides -7 divided by 1x x=-7 then to find Y y=4(-7)+13 y=-28+13 y=-15

jimthompson5910 (jim_thompson5910):

when you have -3x+6=4x+13, you do NOT subtract 3x from both sides. you instead add 3x to both sides

jimthompson5910 (jim_thompson5910):

this is because the initial -3x adds to 3x to get -3x + 3x = 0x = 0 and the x terms go away on the left side

jimthompson5910 (jim_thompson5910):

so if you added 3x to both sides, you go from -3x+6=4x+13 to 6 = 7x + 13

OpenStudy (anonymous):

okay so I have 6=7x+13 do I add 13 or subtract 13?

jimthompson5910 (jim_thompson5910):

you want to undo the +13 on the right side so you subtract 13 from both sides.

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):